madysongodwin:

a gardener wants to make a rectangular enclosure using a wall as one side and 160m of fencing for the other three sides, as shown. For what values of x is the area of the enclosure more than 3000m^2 but less than 3150m^2?

darkknight:

since this is a rectangle, use x as the width and y as the length okay? now assume that the wall that is on one side is y (length) So we know that 160 = y+2x because of the opposite side of the wall (y) and the 2 other sides are the other 2 sides

darkknight:

wait... does it tell you which side is supposed to be x?

darkknight:

length or width?

madysongodwin:

left

madysongodwin:

and length

darkknight:

oh ok, so ignore what i said above. 160 = x +2y and x is the length of the wall

madysongodwin:

okay thank you

madysongodwin:

i have another question if you have time to answer it

darkknight:

we rnt done yet

madysongodwin:

oh okay

darkknight:

so we are finding area so area is xy So we are trying to find a x value that would put the total area between 3000 and 3150. You can really just use guess and check because it isn't asking for an exact value

madysongodwin:

thank you so much :)

madysongodwin:

marco says that the solutions of the equation x^2_6x+13=0 are x=-1 and x=-5. Is marco correct? If so, verify his solutions by solving the equation for x over the set of complex numbers. If not, explain his error and provide the correct solutions

madysongodwin:

can you help me with this?

darkknight:

to get you started if the value of x is 60 then y is 50 puts your area at 3000

Eiwoh2:

New thread pls

darkknight:

and if x is 70 then y is 45 and that puts it at 3150

darkknight:

so have an x value between 60 and 70

madysongodwin:

i got x=30 and y=100 .. 2(30)+ 100=160 30x100=300

darkknight:

read what I wrote above, what you did is not correct. :( But I explained why x should be between 60 and 70 so if u have questions u can ask

darkknight:

I am the darkknight

madysongodwin:

can u answer marco says that the solutions of the equation x^2_6x+13=0 are x=-1 and x=-5. Is marco correct? If so, verify his solutions by solving the equation for x over the set of complex numbers. If not, explain his error and provide the correct solutions

darkknight:

new post

darkknight:

close this one and make a new one

madysongodwin:

ok